There is a total of $40 in dimes and nickels. there is a total of 600 coins. how many are dimes and are many aare nickels.
Answer by checkley77(12844)

D+N=600
D=600-N
10D+5N=4000
10(600-N)+5N=4000
6,000-10N+5N=4000
-5N=4,000-6,000
-5N=-2,000
N=-2,000/-5
N=400 NUMBER OF NICKELS.
600-400=200 NUMBER OF DIMES.
PROOF:
10*200+5*400=4,000
2,000+2,000=4,000
4,000=4,000